### v1

# Extracting Interpretable Logic Rules from Graph Neural Networks

Abstract
Graph neural networks (GNNs) operate over both input feature spaces and combinatorial graph structures, making it challenging to understand the rationale behind their predictions. As GNNs gain widespread popularity and demonstrate success across various domains, such as drug discovery, studying their interpretability has become a critical task. To address this, many explainability methods have been proposed, with recent efforts shifting from instance-specific explanations to global concept-based explainability. However, these approaches face several limitations, such as relying on predefined concepts and explaining only a limited set of patterns. To address this, we propose a novel framework, LogicXGNN, for extracting interpretable logic rules from GNNs. LogicXGNN is model-agnostic, efficient, and data-driven, eliminating the need for predefined concepts. More importantly, it can serve as a rule-based classifier and even outperform the original neural models. Its interpretability facilitates knowledge discovery, as demonstrated by its ability to extract detailed and accurate chemistry knowledge that is often overlooked by existing methods. Another key advantage of LogicXGNN is its ability to generate new graph instances in a controlled and transparent manner, offering significant potential for applications such as drug design. We empirically demonstrate these merits through experiments on real-world datasets such as MUTAG and BBBP.
Machine Learning, ICML
1 Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for modeling and analyzing graph-structured data, achieving remarkable performance across diverse domains, including drug discovery (Xiong et al., 2021; Liu et al., 2022; Sun et al., 2020), fraud detection (Rao et al., 2021), and recommender systems (Chen et al., 2022). Despite their success, GNNs share the black-box nature inherent to the neural network family, which poses challenges to their further development in high-reliability applications such as healthcare (Amann et al., 2020; Bussmann et al., 2021).
To this end, several explainability methods have been developed to uncover the inner decision-making mechanisms of GNNs. However, most of these methods are limited to providing local explanations tailored to specific input instances or rely on interpretations based on input feature attributions (Pope et al., 2019; Ying et al., 2019; Vu & Thai, 2020; Lucic et al., 2022; Tan et al., 2022). Another line of research focuses on global explanations that describe the overall behavior of models (Xuanyuan et al., 2023; Azzolin et al., 2023). These approaches offer more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. However, they have limitations, such as relying on predefined concepts and generating rules that explain only a limited set of patterns within each class without effectively distinguishing between these classes.
To this end, we propose LogicXGNN, a novel framework that extracts interpretable logic rules to explain the internal reasoning process of GNNs while maintaining classification accuracy comparable to the original model. LogicXGNN ensures interpretability by grounding the hidden predicates of the rules into the input space using decision trees, which are both computationally efficient and data-driven. To the best of our knowledge, LogicXGNN is the first interpretable, rule-based functional equivalent of GNNs. Additionally, it can function as a generative model, with notable potential in fields such as drug design and knowledge discovery. We demonstrate the effectiveness of our approach through experiments on real-world datasets, including the MUTAG and BBBP datasets. In summary, we make the following contributions:
  • We propose LogicXGNN, a novel post-hoc approach for extracting interpretable logic rules from graph neural networks. Our method incurs minimal computational overhead, comparable to training decision trees, and is model-agnostic, compatible with a wide range of GNN architectures.
  • LogicXGNN not only generates discriminative rules that serve as a rule-based counterpart to the original model, achieving comparable classification accuracy, but also facilitates knowledge discovery.
  • As LogicXGNN makes decision-making in GNNs transparent with descriptive rules, it can be used as a generative model for creating graph instances, holding significant potential in fields such as drug design.
2 Preliminary
2.1 Graph Neural Networks
Consider a graph $G=(V_{G},E_{G})$, where $V_{G}$ represents the set of nodes and $E_{G}$ represents the set of edges. For a collection of graphs $\mathcal{G}$, let $\mathcal{V}$ and $\mathcal{E}$ denote the sets of vertices and edges across all graphs in $\mathcal{G}$, respectively, with $|\mathcal{V}|=n$. Each node is associated with a $d_{0}$-dimensional feature vector, and the input features for all nodes are represented by a matrix $\mathbf{X}\in\mathbb{R}^{n\times d_{0}}$. An adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ is defined such that $\mathbf{A}_{ij}=1$ if an edge $(i,j)\in\mathcal{E}$ exists, and $\mathbf{A}_{ij}=0$ otherwise. A graph neural network (GNN) model $\mathcal{M}$ learns to embed each node $v\in\mathcal{V}$ into a low-dimensional space $\mathbf{h}_{v}\in\mathbb{R}^{d_{L}}$ through an iterative message-passing mechanism over the $L$ number of layers. At each layer $l$, the node embedding is updated as follows:
$\mathbf{h}_{v}^{l+1}=\text{UPD}\left(\mathbf{h}_{v}^{l},\text{AGG}\left(\left\{\text{MSG}(\mathbf{h}_{v}^{l},\mathbf{h}_{u}^{l})\mid\mathbf{A}_{uv}=1\right\}\right)\right),$
where $\mathbf{h}_{v}^{0}=\mathbf{X}_{v}$ is the initial feature vector of node $v$, and $\mathbf{h}_{v}^{l}$ represents the final node embedding at the final layer $L$. The update function UPD, aggregation operation AGG, and message function MSG define the architecture of a GNN. For instance, Graph Convolutional Networks (GCN) (Kipf & Welling, 2017) use an identity message function, mean aggregation, and a weighted update. Non-linearities are incorporated into the model through activation functions, typically the Rectified Linear Unit (ReLU), applied as: $\mathbf{h}_{v}^{l}=\text{ReLU}(\mathbf{h}_{v}^{l})$. In essence, the GNN $\mathcal{M}$ aggregates information from both the feature space and the topological structure of $G$ to compute deep hidden embeddings for nodes. These embeddings are then optimized for specific downstream tasks, such as node or graph classification.
Node Classification
We assume a label function $f:\mathcal{V}\rightarrow\{1,\dots,C\}$ that assigns one of $C$ classes to each node in $\mathcal{V}$. To approximate $f$, we define a GNN model $\mathcal{M}$ for node classification by passing the final node embeddings $\mathbf{h}_{v}^{(L)}$ through a softmax function for each node $v$.
Graph Classification
Similarly, suppose we have a set of graphs $\mathcal{G}$ and a label function $f:\mathcal{G}\rightarrow\{1,\dots,C\}$ that assigns one of $C$ classes to each graph in $\mathcal{G}$. To approximate $f$, we define a GNN model $\mathcal{M}$ for graph classification by passing the graph embeddings $\mathbf{h}_{G}^{L}$ to a fully connected layer followed by a softmax function. Here, the graph embeddings are commonly computed by taking the mean of all node embeddings in the graph $\mathbf{h}_{G}^{L}:=\text{mean}(\mathbf{h}_{v}^{L}\mid v\in V_{G})$ through the operation global_mean_pooling.
2.2 First-Order Logic Rules for GNN Interpretability
First-order logic (FOL), also known as predicate logic, is highly interpretable to humans, making it an excellent tool for explaining the behaviour of neural networks (Zhang et al., 2021). In this paper, our proposed framework, LogicXGNN, aims to elucidate the inner decision-making process of a GNN $\mathcal{M}$ using a Disjunctive Normal Form (DNF) formula $\phi_{M}$. The formula $\phi_{M}$ is a logical expression that can be described as a disjunction of conjunctions (OR of ANDs) over a set of predicates $P$, where each $p_{j}$ represents a logical condition or property defined on the graph structure $\mathbf{A}$ and input features $\mathbf{X}$. Importantly, $\phi_{M}$ incorporates the universal quantifier ($\forall$), providing a global explanation that is specific to a class of instances. This makes $\phi_{M}$ a rule-based model that is functionally equivalent to $\mathcal{M}$ a significant advantage not typically offered by other explanation work, such as attribute-based approaches. To provide a concrete example of $\phi_{M}$, suppose the model $M$ is trained to determine whether a molecule $G$ is soluble. In this case, we may extract the following logical rules $\phi_{M}$ from $M$:
$\forall G,\,(p_{1}(G)\land p_{2}(G))\lor(p_{3}(G)\land p_{4}(G))\Rightarrow\text{label}(G,s),$
where $p_{1}(G)$ represents “The molecule contains a hydroxyl group (-OH)”, $p_{2}(G)$ represents “The molecule has a ring structure”, $p_{3}(G)$ represents “The molecule contains a carbonyl group (C=O)”, $p_{4}(G)$ represents “The molecule has a high degree of branching”, and $\text{label}(G,s)$ is a predicate indicating that graph $G$ is assigned the class label $s$ (“soluble”). Similarly, for node classification tasks, predicates $\{p_{j}(v)\}$ encode logical conditions and properties on nodes’ neighbourhoods and features.
While looks promising, extracting such a DNF formula $\phi_{M}$ from the original GNN $\mathcal{M}$ is indeed challenging. More specifically, we need to address the following key questions:
  1. 1.
    How can we identify a set of predicates $P$ that are both interpretable and critical for classification?
  2. 2.
    How can we determine the logical structure of $\phi_{M}$ that not only explains data from a specific class but also effectively rejects data from other classes?
  3. 3.
    Can we design an approach that is both efficient (with minimal computational overhead) and generalizable to different tasks and model architectures?
We show how LogicXGNN addresses these challenges in Section 3. Our key observation is that neural networks, including GNNs, exhibit class-specific activation patterns in their hidden representational spaces, capturing diverse instances within the same class (Yosinski et al., 2015; Gopinath et al., 2019; Geng et al., 2023, 2024). The central idea of LogicXGNN is to leverage these hidden patterns as predicates and then identify logical structures in a data-driven manner, ultimately grounding the predicates into the input space. This design choice aligns with the spirit of a recent work, NeuroLogic (Geng et al., 2025), which shows promising results in explaining fully connected networks and convolutional neural networks. Interested readers can refer to the original paper for further details.
3 The LogicXGNN Framework
This section provides a detailed discussion of our methodology and design choices in LogicXGNN. For ease of discussion, we divide LogicXGNN into three critical sub-problems. Figure LABEL:fig:overview illustrates an overview of the LogicXGNN framework.
3.1 Identifying Hidden Predicates $P$ for $\phi_{M}$
We start by discussing the identification of hidden predicates for graph classification tasks. In fact, we define hidden predicates at the node level, so it naturally works for node classification tasks, as will be clearly explained later.
As previously mentioned, the desired predicates $P$ should capture commonly shared patterns in both graph structures $\mathbf{A}$ and hidden embeddings $\mathbf{h}^{L}$ across a set of instances in the context of GNNs. While graph structure information can be encoded into hidden embeddings, it often becomes indistinguishable due to oversmoothing during the message-passing process. To mitigate this, we explicitly model common patterns in graph structures.
After $L$ layers of message passing, the receptive field of a node $v$ corresponds to a subgraph that includes the node itself and its $1,\dots,L$-hop neighborhoods. Intuitively, nodes with isomorphic receptive fields tend to exhibit similar properties and may even belong to the same class. Similarly, graph instances sharing multiple isomorphic subgraphs often display related characteristics. To exploit this, we use these subgraphs—nodes’ neighborhoods—to represent structural patterns and use graph hashing to compare and store these patterns efficiently. Formally, the computation of a structural pattern contributed by a certain node $v$ is given by the following function:
$\text{Pattern}_{\text{struct}}(v)=\text{Hash}\big{(}\text{ReceptiveField}(v,\mathbf{A},L)\big{)}.$
Next, we discuss common patterns in the hidden embeddings. During the training of GNNs for classification tasks, the hidden embeddings are optimized to differentiate between classes. Empirically, we find that a small subset of specific dimensions in the final-layer embeddings $\mathbf{h}_{G}^{L}$ is sufficient to distinguish instances from different classes when using appropriate thresholds, often achieving similar accuracy to the original neural networks. Similar observations have been reported in (Geng et al., 2025).
In this work, we apply the decision tree algorithm to the collection of final-layer graph embeddings of training data to identify a set of the most informative dimensions $K$ along with their corresponding thresholds $T$. Formally, this is expressed as:
$\text{DecisionTree}(\{\mathbf{h}_{G}^{L}\mid G\in\mathcal{G}\},\mathbf{Y})\rightarrow(K,T)$
where $\mathbf{Y}$ represents the label vector. We then leverage this information to construct embedding patterns at the node level, aligning with the definition of structural patterns. Recall that $\mathbf{h}_{G}^{L}:=\text{mean}(\mathbf{h}_{v}^{L}\mid v\in V_{G})$, so we broadcast $K$ and $T$ to each node embedding $\mathbf{h}_{v}^{L}$. For node classification tasks, since $K$ and $T$ are already computed at the node level, broadcasting is unnecessary. Then, for an input node $v$, its embedding value $\mathbf{h}_{v}^{L}$ at each informative dimension $k\in K$ is compared against the corresponding threshold $T_{k}$. The result is then abstracted into binary states: 1 (activation) if the condition is met, and 0 (deactivation) otherwise. Formally, we have:
$\mathcal{I}_{k}({\mathbf{h}_{v}^{L}})=\begin{cases}1&\text{if }\mathbf{h}_{v}^{L}[k]\geq T_{k}\\ 0&\text{if }\mathbf{h}_{v}^{L}[k]<T_{k}\end{cases}$
In summary, for both node and graph tasks, the embedding pattern contributed by a given node $v$ can be computed using the following function:
$\text{Pattern}_{\text{emb}}(v)=\left[\mathcal{I}_{1}(\mathbf{h}_{v}^{L}),\mathcal{I}_{2}(\mathbf{h}_{v}^{L}),\dots,\mathcal{I}_{K}(\mathbf{h}_{v}^{L})\right]$
Putting it together, we define the predicate function as $f(v)=(\text{Pattern}_{\text{struct}}(v),\text{Pattern}_{\text{emb}}(v))$. To identify the set of hidden predicates, we iterate over each node $v\in\mathcal{V}$ in the training set, collect all $f(v)$, and transform them into a set $P$. In addition, when a node $v$ is evaluated against a predicate $p_{j}$, the evaluation $p_{j}(v)$ is true only if both the structural and embedding patterns from $f(v)$ match the predicate. To extend the applicability of a predicate to a graph instance $G$, we override its definition as follows:
$p_{i}(G)=1\text{ if }\exists v\in V_{G},\,p_{i}(v)=1$
$p_{i}(G)=0\text{ if }\forall v\in V_{G},\,p_{i}(v)=0$
To better illustrate the process of identifying hidden predicates, we present a simple example in Figure LABEL:fig:graph_a. This scenario involves a binary graph classification task, a common setup in GNN applications. In this example, we have five input graphs, with each node characterized by two attributes: degree and type. The types, “A” and “B”, are encoded as 0 and 1, respectively. A GNN with a single message-passing layer is applied, generating a 2-dimensional embedding for each node (i.e., $d_{L}=2$). As only one message-passing layer is used, structural patterns are extracted based on the nodes and their first-order neighbors.
Using decision trees, we identify the most informative dimension $k=1$, and its corresponding threshold $t=0.18$ from the graph embeddings. This threshold is then applied to the node embeddings to compute embedding patterns. As a result, six predicates are derived. Notably, $p_{5}$ (“83c89e”, 1) and $p_{6}$ (“83c89e”, 0) exhibit isomorphic structures—represented by identical hash strings—but differ in their embedding activations. We conclude that these predicates are: 1) fundamental building blocks of graph instances, as they are iterated over all structural patterns in the training data, and 2) critical for classification, since the embedding patterns align with the results of the decision tree.
3.2 Determining the Logical Structure of $\phi_{M}$
In this subsection, we aim to construct global logical rules $\phi_{M}$ based on hidden predicates $P$ for each class, which serve the same functionality as the original GNN $M$. To achieve this, we adopt a data-driven approach. We process all training instances from class $c\in C$ that are correctly predicted by $M$, evaluating them against the predicates $P$ and recording their respective activation patterns. The results are stored in a binary matrix $\Phi_{c}$ for each class $c$, where the columns correspond to the predicates in $P$, and the rows represent the training instances correctly classified as $c$ by $M$. Specifically, an entry $\Phi_{c}[i,j]=1$ denotes that the $j$-th instance exhibits the $i$-th predicate, while $\Phi_{c}[i,j]=0$ indicates otherwise, as illustrated in Figure LABEL:fig:graph_b.
From a logical structure perspective, each row in $\Phi_{c}$ represents a logical rule that describes an instance of class $c$, expressed in conjunctive form using hidden predicates. For instance, in the simple binary classification task introduced earlier, $G_{1}$ corresponds to the column $(1,1,0,1,0,0)$, which can be represented as $p_{1}\land p_{2}\land p_{4}$. To derive the global descriptive rules for class $c$, denoted as $\bar{\phi}_{M}^{c}$, we take the disjunction (OR) of all distinct conjunctive forms. Thus, the global rule for our GNN $M$ in the simple binary classification task, denoted as $\bar{\phi}_{M}$, can be expressed as follows:
$\forall G,\text{ }$ $\left(p_{1}\land p_{2}\land p_{4}\right)\lor\left(p_{1}\land p_{5}\right)\Rightarrow\textit{label}(G,0),$
$\forall G,\text{ }$ $\left(p_{1}\land p_{2}\land p_{3}\right)\lor\left(p_{1}\land p_{3}\land p_{4}\land p_{6}\right)\Rightarrow\textit{label}(G,1).$
Here, we omit $G$ in $p_{j}(G)$ when the context is clear. In addition to the descriptive rules $\bar{\phi}_{M}^{c}$, we also record and learn the recurring connectivity patterns of predicates for each conjunctive form. For instance, in the case of $(p_{1}\land p_{2}\land p_{4})$, both $p_{1}$ and $p_{2}$ are connected to $p_{4}$, which can be represented using an adjacency matrix. Multiple connectivity patterns may exist, and these can be learned in a data-driven manner, similar to how $\bar{\phi}_{M}$ is learned. We denote the collection of connectivity patterns as $\psi_{M}$, which is useful for graph generation and motif-level rule grounding.
To derive an even more compact set of rules that effectively distinguish between classes, we input $\Phi$ and $Y$ into a decision tree. The decision tree then identifies the most distinctive and discriminative rules, denoted as $\hat{\phi}_{M}$. For example, in our simple GNN case, the decision tree generates the following discriminative rules:
$\forall G,\ \neg p_{3}(G)\Rightarrow\textit{label}(G,0),\quad\forall G,\ p_{3}(G)\Rightarrow\textit{label}(G,1).$
Note that the discriminative rules $\hat{\phi}_{M}$ do not fully replace the descriptive rules $\bar{\phi}_{M}$, as the latter remain valuable for generating graph instances—a task that discriminative rules alone cannot accomplish. Both $\hat{\phi}_{M}$ and $\bar{\phi}_{M}$ constitute the rules extracted from model $M$, collectively denoted as $\phi_{M}$.
3.3 Grounding $\phi_{M}$ into the Input Space
The next challenge is interpreting $\phi_{M}$. To address this, we focus on grounding its building blocks—predicates $P$—into the input space $\mathbf{X}$, thereby bridging the abstract logic with tangible input features. Recall that a predicate $p_{j}$ encodes a subgraph centered on a node $v$, encompassing its $1,\dots,L$-hop neighborhoods. Each neighbour node is also mapped to some predicate by the predicate function $f$. Thus, we can leverage the input features of the node $v$ along with its neighbour predicates to learn rules that relate the predicate $p_{j}$ to the input space. To achieve this, we define the input feature for the subgraph centered at $v$ as $\mathbf{Z}_{v,L}$, which is constructed by concatenating the information of nodes at each neighborhood level $l$:
$\mathbf{Z}_{v,L}=\text{CONCAT}\left(\mathbf{X}_{v},\text{Encode}\left(\{f(u)\mid u\in\mathcal{N}^{(1)}(v)\}\right),\right.$
$\left.\dots,\text{Encode}\left(\{f(u)\mid u\in\mathcal{N}^{(L)}(v)\}\right)\right)$
where $\mathcal{N}^{(l)}(v)$ represents the set of $l$-hop neighbors of $v$, and Encode transforms the collection of predicates into a frequency-based encoding, serving as an order-invariant aggregation operator applied to the predicates of the $l$-hop neighbors. For example, as illustrated in Figure LABEL:fig:graph_c, the subgraph input for node 1 is $\mathbf{Z}_{1,1}=(1,A,p_{4})$, which indicates that node 1 has a degree of 1 and a type of “A”. It also connects to predicate $p_{4}$, representing its sole neighbor, node 2. Similarly, the subgraph input for node 2 is $\mathbf{Z}_{2,1}=(3,B,p_{1},p_{1},p_{2})$. Here, we omit the encoding in $\mathbf{Z}$ for demonstration purposes.
Since model-agnosticism is a key requirement for LogicXGNN, we approximate the message-passing layers using interpretable models. In this work, we utilize decision trees for this purpose. Specifically, decision trees are employed to generate input-level rules that distinguish predicates with isomorphic subgraphs but different embedding patterns. For this task, the training data for each predicate label $j$ is the set of subgraph inputs of the nodes $v$ that activate $p_{j}$, represented as $\{\mathbf{Z}_{v,L}\mid p_{j}(v)=1\}$. For example, the training data for $p_{1}$ (identified as (“cde85e”, 0)) is $\{\mathbf{Z}_{1,1},\mathbf{Z}_{3,1},\dots,\mathbf{Z}_{22,1}\}$, while the training data for $p_{2}$ (identified as (“cde85e”, 1)) is $\{\mathbf{Z}_{4,1},\mathbf{Z}_{11,1},\mathbf{Z}_{20,1}\}$. The decision tree generates input-level rules such as $z[1]\leq 0.5$ for $p_{1}$, and otherwise for $p_{2}$. Given the prior knowledge that $z[1]$ represents the node’s type, we recognize that $p_{1}$ encodes whether the node is of type “A”, and $p_{2}$ is true when the node is of type “B”.
On the other hand, for predicates that do not have an isomorphic counterpart, we simply extract the representative features - this can be selected based on expert knowledge or those with the smallest variance — to explain the corresponding predicates. For instance, for $p_{3}$, supported by $\{\mathbf{Z}_{18,1},\mathbf{Z}_{19,1},\mathbf{Z}_{23,1},\mathbf{Z}_{24,1}\}$, the feature with the smallest variance is dimension 0. This indicates that $p_{3}$ encodes the fact that the node has a degree of 2. Combining the discriminative rules derived earlier, we yield the following interpretable logic rules $\hat{\phi}_{M}$ from the model $M$:
$\forall G,\,\forall v\in V_{G}\,\left(\text{degree}(v)\neq 2\right)$ $\Rightarrow\textit{label}(G,0),$
$\forall G,\,\exists v\in V_{G}\,\left(\text{degree}(v)=2\right)$ $\Rightarrow\textit{label}(G,1).$
Sometimes, it is necessary to include neighbor predicates for effective grounding. For instance, $p_{4}$, supported by $\{\mathbf{Z}_{2,1},\mathbf{Z}_{12,1}\}$, encodes the presence of predicate $p_{2}$ as a neighbor. Since $p_{2}$ encodes type “B,” $p_{4}$ can be interpreted as having a neighbor node of type “B”. By incorporating connectivity patterns $\psi_{M}$, we can generate motif-level rules rather than node-level through the grounding of adjacent predicates. For example, in our simple GNN case, as shown in Figure LABEL:fig:graph_c, $\psi_{M}$ learns that two instances of $p_{3}$ are connected and share connections with the same other predicates. Given that these two nodes corresponding to $p_{3}$ have a degree of 2, we can directly infer the following rule:
$\forall G,\,\neg\text{has\_cycle}(G)\Rightarrow\textit{label}(G,0),$
$\forall G,\,\text{has\_cycle}(G)\Rightarrow\textit{label}(G,1)$
which aligns more closely with human observation, despite the node-level rules are also correct.
3.4 Extensions
Inference with LogicXGNN
Unlike a graph neural network $M$, which classifies based on the highest logit, its rule-based equivalent $\phi_{M}$ classifies an instance as belonging to a class if two conditions are satisfied simultaneously: 1) The instance matches that class’s (discriminative) rules, and 2) it does not match any other class’s rules. On the other hand, if an instance matches the rules of multiple classes, the model reports the conditions instead of making a definitive choice. This approach is safer, as $\phi_{M}$ only makes decisions when confident and provides the reasoning—matched rules—for human inspection in ambiguous cases.
LogicXGNN as Generative Model
As LogicXGNN makes each decision-making step in GNNs transparent with descriptive rules, we can leverage LogicXGNN as a generative model for creating graph instances in a controlled manner. First, we sample a conjunctive form from $\bar{\phi}_{M}$, then combine it with the connectivity pattern $\psi_{M}$ to form a graph consisting of predicates, and finally, assign node features based on the grounding rules associated with the corresponding predicates. We provide more details in Appendix A.
Table 1: Results of the cardinality of the extracted rules $\phi_{M}$, along with a comparison of the train and test accuracies fo $\hat{\phi}_{M}$, the original GNN $M$, and GLGExplainer. The columns labeled $|P|$, $|\bar{\phi}_{M}^{0}|$, $|\bar{\phi}_{M}^{1}|$, $|\hat{\phi}_{M}^{0}|$, and $|\hat{\phi}_{M}^{1}|$ report the number of predicates, as well as the number of conjunction clauses in the descriptive rules $\bar{\phi}_{M}$ and discriminative rules $\hat{\phi}_{M}$ for classes 0 and 1, respectively.
The cardinality of $\phi_{M}$ Train accuracy Test accuracy
Dataset $|P|$ $|\bar{\phi}_{M}^{0}|$ $|\bar{\phi}_{M}^{1}|$ $|\hat{\phi}_{M}^{0}|$ $|\hat{\phi}_{M}^{1}|$
$M$
$\hat{\phi}_{M}$
GLG.
$M$
$\hat{\phi}_{M}$
GLG.
HIN 2293 426 425 146 62 89.12 93.83 50.04 86.93 85.23 49.99
BBBP 254 146 1044 20 22 81.18 87.37 45.18 81.13 81.37 42.90
Mutagenicity 314 827 1015 198 167 78.12 84.03 56.78 76.04 76.50 54.26
Figure 3: Selected knowledge discovered by LogicXGNN. The rows display results for the BBBP and MUTAG datasets, respectively.
4 Evaluation
In this section, we conduct experimental evaluations on real-world datasets to address the following research questions:
  1. 1.
    How effective is $\phi_{M}$ as a classification tool? Does it have the potential to replace the original GNN $M$?
  2. 2.
    What knowledge can we derive from the underlying benchmark using $\phi_{M}$?
  3. 3.
    How effective is $\phi_{M}$ as a generative model, and what are its advantages?
Benchmarks and Baselines
We use the following graph classification benchmarks: (1) The Mutagenicity (MUTAG) dataset consists of 4337 molecule graphs, each labeled based on their mutagenic effect on the Gram-negative bacterium Salmonella typhimurium (Debnath et al., 1991). (2) The BBBP dataset consists of 2039 molecule graphs, each labeled based on their ability to penetrate the blood-brain barrier (Wu et al., 2018). (3) The Hospital Interaction Network (HIN) dataset comprises 1760 graphs representing face-to-face interactions between doctors and nurses. The task is to classify whether the ego in each ego network is a doctor or a nurse (Azzolin et al., 2023). For the baseline approaches, we employ a generation-based method, XGNN (Yuan et al., 2020), as well as two concept-based methods, GCNeuron (Xuanyuan et al., 2023) and GLGExplainer (Azzolin et al., 2023). All these methods are considered state-of-the-art.
Setup
All experiments are conducted on an Ubuntu 22.04 LTS machine with 32 GB of RAM and a 2.7 GHz processor. We use a 2-layer GNN with a hidden dimension of 32 for all datasets. To demonstrate the model-agnostic nature of LogicXGNN, we employ graph convolutional networks for BBBP and HIN, and GraphSAGE (Hamilton et al., 2017) for MUTAG. All datasets are split into train-test sets with an 8:2 ratio. Additionally, we use the CART algorithm for the decision trees in LogicXGNN.
Figure 4: Selected examples generated by descriptive rules $\bar{\phi}_{M}$ and explanation graphs generated by XGNN. The rows display results for BBBP, MUTAG, and HIN, respectively. Each cell under LogicXGNN includes the original graphs, their descriptive rules, and a newly generated instance based on these rules.
4.1 How effective is $\phi_{M}$ as a classification tool?
We report the cardinality of $\phi_{M}$, the rules extracted by LogicXGNN from the original GNN $M$, and compare its classification accuracy on both the train and test sets with $M$ and baseline approaches in Table 1.
Notably, our rule-based model $\hat{\phi}_{M}$ outperforms its neural counterpart $M$ on the BBBP and MUTAG datasets, despite being derived from it. While $\hat{\phi}_{M}$ shows a slight tendency to overfit the training data, it generalizes quite well on the test data, achieving nearly identical performance to the original model. For example, on the BBBP dataset, $\hat{\phi}_{M}$ achieves an accuracy of $81.37\%$ on the test set, marginally outperforming the $81.13\%$ accuracy of the original GNN $M$. This suggests that $\hat{\phi}_{M}$ successfully captures the underlying decision-making logic of $M$, demonstrating its potential to replace neural models in many settings where high fidelity and interpretability are critical. Having a white-box model like $\hat{\phi}_{M}$ offers many significant advantages. For instance, unlike black-box models, it provides reasoning in the form of interpretable logic rules rather than merely producing predictions. More importantly, it enables users to inspect where the model may make errors, allowing for targeted improvements by incorporating human prior knowledge to further enhance the model’s performance. For instance, the number of carbon rings influences mutagenicity, yet this information is not preserved in the global_mean_pooling operation of GNNs. In contrast, our rule-based model can explicitly incorporate this factor with a simple patch. We leave this direction for future work.
As for the baselines, GCneuron lacks discriminative power, as it generates explanations based solely on the input class. Specifically, providing different target classes for the same input yields entirely different explanations. On the other hand, GLGExplainer underestimates the rich combinatorial complexity of graphs and tends to produce overly simplistic rules that fail to capture the variety of data patterns. For instance, GLGExplainer uses only two predicates, $p_{0}$ and $p_{1}$, to represent explanations for classes 0 and 1 in the Mutagenicity dataset. However, to accurately distinguish between these classes, our approach generates disjunctive normal forms (DNF) of sizes 198 and 167 for classes 0 and 1, respectively. As a result, GLGExplainer’s performance is similar to a random classifier.
4.2 What knowledge can we derive using $\phi_{M}$?
We ground the discriminative rules $\hat{\phi}_{M}$ and present the selective knowledge extracted from $M$ for the BBBP and MUTAG datasets in Figure 3. These findings are accurate from the perspectives of chemistry and molecular science, and they effectively explain the classification results. For instance, we recognize that molecules containing oxygen-rich groups are less likely to cross the Blood-Brain Barrier (class 0). This is due to their increased hydrophilicity, reduced lipophilicity, and higher likelihood of being recognized by efflux transporters (Pardridge, 2005).
Regarding mutagenicity, we find that molecules containing a methyl group (-$\text{-CH}_{3}-$) are non-mutagenic (class 1). This is because the methyl group is generally less reactive and does not readily participate in chemical reactions that could lead to DNA damage or mutagenesis (Jin & Robertson, 2012). Additionally, we find that the combination of an aromatic or heteroaromatic ring with a nitro group is crucial for mutagenic activity (class 0), consistent with the findings of the original paper (Debnath et al., 1991). In contrast, existing approaches like GCNeuron and GLGExplainer treat nitro groups and carbon rings as isolated entities, overlooking the critical synergy between them.
4.3 How effective is $\phi_{M}$ as a generative model?
We present selected examples generated by descriptive rules $\bar{\phi}_{M}$ and explanation graphs generated by the baseline approach, XGNN, in Figure 5. As the descriptive rules make each decision-making step in GNNs transparent, we can accurately model the underlying data distribution, enabling the generation of new instances in a fully controlled manner. Specifically, we can select certain conjugate clauses from the descriptive rules, particular connection patterns, and preferences in predicate grounding during the generation process. For instance, we select the conjugate clause $p_{41}\land p_{6}\land p_{15}$ with the following connection pattern: $p_{15}$ is connected to $p_{61}$, $p_{61}$ is connected to $p_{6},p_{6},p_{15}$, and $p_{6}$ is connected to $p_{6},p_{61}$. Given the grounding rule that $p_{41},p_{6},$ and $p_{15}$ all encode motifs composed of carbon atoms, we successfully reconstruct the original molecule.
In contrast, XGNN learns a reinforcement learning agent that is inherently black-boxed, generating graphs that do not align with the actual data distribution. For example, it generates bipartite structures that lack molecular relevance, whereas our approach learns the underlying rules and adheres to them during the generation process. As a result, $\bar{\phi}_{M}$ can not only reconstruct original instances but also generate new ones that preserve similar structures and key properties.
We envision adding more flexibility to the generation process, such as randomly combining conjugate clauses from descriptive rules, which could lead to even more interesting results. This approach may enable newly generated molecular instances to inherit key properties from multiple known molecules. In summary, LogicXGNN demonstrates significant potential in fields such as drug design, and we will further explore these possibilities in future work.
5 Related Work
The explainability of GNNs remains a relatively underexplored area compared to other neural networks, such as convolutional neural networks (CNNs). Most existing methods focus on providing local input attribution explanations (Pope et al., 2019; Ying et al., 2019; Luo et al., 2020b; Vu & Thai, 2020; Lucic et al., 2022; Tan et al., 2022), similar to attribute-based approaches like Grad-CAM (Selvaraju et al., 2017) used for explaining CNNs. Another line of research focuses on global explanations that capture the overall behavior of models, which is where our approach, LogicXGNN, also belongs. Global explanation methods can be broadly divided into generation-based and concept-based approaches. Generation-based methods, such as XGNN (Yuan et al., 2020), use reinforcement learning agents to generate graph instances that maximize specific model predictions. Similarly, GNNInterpreter (Wang & Shen, 2023) learns a probabilistic generative graph distribution to produce graph patterns that serve as explanations. These approaches are black-boxed and provide only a limited number of explanation patterns for certain classes, whereas LogicXGNN is transparent and can generate any desired instance in a controlled manner, making it highly applicable to fields like drug design. On the other hand, concept-based approaches aim to provide more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. For example, GCneuron (Xuanyuan et al., 2023) identifies predefined concepts, formulated as logical combinations of node degrees and neighborhood properties, associated with specific neurons. Similarly, GLGExplainer (Azzolin et al., 2023) builds on local explanations from PGExplainer (Luo et al., 2020a), maps them to learned concepts and derives logic formulas from these concepts to serve as class explanations. In contrast, LogicXGNN offers a simpler and more intuitive approach, relying solely on decision tree computations rather than complex local explanation methods. It is data-driven, eliminating the need for predefined concepts. Most importantly, while these concept-based approaches typically generate descriptive rules for individual classes without distinguishing between them, LogicXGNN excels in creating discriminative rules that differentiate classes and matches the original model’s accuracy.
6 Conclusion and Future Work
In this work, we present LogicXGNN, a novel framework for extracting interpretable logic from GNNs. LogicXGNN is architecture-agnostic and focuses on extracting global rules rather than instance-level explanations. Unlike existing global methods, it generates discriminative rules that effectively differentiate between classes while achieving accuracy comparable to the original model. Additionally, LogicXGNN can function as a generative model for creating graph instances, demonstrating considerable potential in applications such as drug design. For future work, we identify two key directions to extend LogicXGNN: 1) Enhancing Rule-based Models with Program Synthesis: Since LogicXGNN translates GNN decision-making into logic/program space, we aim to leverage program synthesis techniques and human prior knowledge to further improve the performance of our rule models; 2) Interpretable Approaches for Regression Tasks: Current methodologies of LogicXGNN are tailored for classification tasks. Extending the framework to handle regression problems, such as predicting molecular properties and 3D structures, remains an exciting challenge for future exploration.
Impact Statement
The interpretability of Graph Neural Networks (GNNs) is an essential aspect of their application, particularly in fields like drug discovery, healthcare, and safety-critical systems, where understanding the rationale behind model predictions can have significant consequences. LOGICXGNN tackles the challenge of interpretability by extracting human-readable logic rules from GNNs without relying on predefined concepts, offering a data-driven approach that makes the decision-making process more transparent. This not only enhances our understanding of GNN predictions but also facilitates the discovery of valuable knowledge that may otherwise be overlooked, as demonstrated by experiments on real-world datasets such as MUTAG and BBBP.
One of the key advantages of LOGICXGNN is its potential to replace GNNs as a classifier in scenarios where high fidelity and interpretability are paramount. This is particularly important in settings such as drug design, where model transparency can help prevent harmful or unintended outcomes. Additionally, by enabling the generation of new graph instances in a controlled and transparent manner, LOGICXGNN opens up new possibilities for innovative and reliable applications across various domains.
From an ethical standpoint, LOGICXGNN promotes greater accountability and trust in AI-driven decisions by providing understandable justifications for its predictions. However, it is crucial to acknowledge the potential risks of over-relying on automated systems, particularly in high-stakes contexts. As such, while LOGICXGNN can complement human expertise, it should not replace human oversight, especially when it comes to critical decision-making processes.
Appendix A LogicXGNN as Generative Model
We introduce Algorithm 1 for generating graph instances with LogicXGNN and illustrate the process with a step-by-step real-world example in Figure 5 for clarity.
1
Input: Descriptive rules ($\bar{\phi}_{M}$), connectivity patterns ($\psi_{M}$), predicate grounding rule ($R$)
Output: A graph $G=(V,E)$ with node features $X_{V}$
2
3 Function generateGraph($\bar{\phi}_{M}$, $\psi_{M}$, $R$)
4
      $C\leftarrow\textit{Sample}(\bar{\phi}_{M})$
        /* Sample a conjugate clause $C$/
5
      $Template\leftarrow\psi_{M}(C)$
        /* Extract connectivity patterns regrading $C$/
6
      $E\leftarrow\emptyset$, $V_{\text{dict}}\leftarrow\{\}$
        /* Initialize edges and nodes */
7
8      for $tmp\in Template$ do
[…]
              /* Compute node frequency */
9
[…]
              /* Update node dictionary */
10
[…]
              /* Unpack pattern tuple */
11
[…]
              /* Select a node randomly */
12
[…]
15                   $E\leftarrow E\cup\{(\text{node\_picked},j)\}$
[…]
                    /* Add bidirectional edges */
16
17
18
      $X_{V}\leftarrow R(V_{\text{dict}},E)$
        /* Apply grounding rule once graph generation completes */
19
20      return $(V_{\text{dict}},E),X_{V}$
Algorithm 1 Graph Instance Generation
Figure 5: A step-by-step example.